OpenStudy (anonymous):

1/2=e^-2t, how do you solve for t? (Answer: (ln 2)/2.)

OpenStudy (anonymous):

Take ln on both sides resulting in ln(1/2)=-2t. Next, divide by -2, resulting in -ln(1/2)/2. The rules of logarithms allows us to claim that ln(1/2)=-ln(2) allowing us to say that the answer is ln(2)/2.